Using the formula B=P+Prt
Suppose an account earns 5.5% interest and after two years the balance is $3330. Find the principal that was invested.
So far I have 3,330= P+P(.55)(2) but I can't figure out what i'm doing wrong and my book isn't helping :/

Answer by stanbon(75887)

So far I have 3,330= P+P(.55)(2)
------
3,330 = P(1+1.10)
------
3330 =P(2.1)
------
P = 3330/2.1
------
P = $1585.71
==================
Cheers,
Stan H.
----------